# RetrievalQA: Assessing Adaptive Retrieval-Augmented Generation for Short-form Open-Domain Question Answering

### v1

@@ Abstract
Adaptive retrieval-augmented generation (ARAG) aims to dynamically determine the necessity of retrieval for queries instead of retrieving indiscriminately to enhance the efficiency and relevance of the sourced information. However, previous works largely overlook the evaluation of ARAG approaches, leading to their effectiveness being understudied. This work presents a benchmark, RetrievalQA, comprising 1,271 short-form questions covering new world and long-tail knowledge. The knowledge necessary to answer the questions is absent from LLMs; therefore, external information must be retrieved to answer correctly. This makes RetrievalQA a suitable testbed to evaluate existing ARAG methods. We observe that calibration-based methods heavily rely on threshold tuning, while vanilla prompting is inadequate for guiding LLMs to make reliable retrieval decisions. Based on our findings, we propose Time-Aware Adaptive REtrieval (TA-ARE), a simple yet effective method that helps LLMs assess the necessity of retrieval without calibration or additional training111The dataset and code will be available at https://github.com/hyintell/RetrievalQA.
@@ 1 Introduction
Retrieval-augmented generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023) that augments large language models (LLMs) with retrieval of relevant information has become increasingly popular in knowledge-intensive tasks, including open-domain question-answering (QA) (Zhang et al., 2023; Kasai et al., 2023; Cui et al., 2023). However, standard RAG methods conduct retrieval indiscriminately, irrespective of the input query, which may result in suboptimal task performance and increased inference costs (Gao et al., 2024). On one hand, LLMs encode vast knowledge in parameters through large-scale pre-training, enabling them to effortlessly handle straightforward queries without retrieval (Mallen et al., 2023). On the other hand, the retrieved context may contain noise and irrelevant information, and augmenting noisy context can potentially distract LLMs, thereby impeding task performance (Shi et al., 2023).
@@
To fill this gap, we create RetrievalQA, a short-form QA dataset with 1,271 questions, covering new world and long-tail knowledge and spanning diverse topics. We ensure the knowledge necessary to answer the questions is absent from LLMs. Therefore, LLMs must truthfully decide whether to retrieve to be able to answer the questions correctly. RetrievalQA enables us to evaluate the effectiveness of ARAG approaches, an aspect predominantly overlooked in prior studies and recent RAG evaluation systems (Chen et al., 2023; Saad-Falcon et al., 2023; Es et al., 2023), which focus only on task performance, the relevance of retrieval context or the faithfulness of answers.
@@
| Category | Data Source | Question | Answer | |||||
| New world knowledge |
Former Texas Rep. Will Hurd | |||||
Mean Girls | ||||||
| Long-tail knowledge |
8:00 PM | |||||
What is Henry Feilden’s occupation? | politician | ||||||
fieldfare |
@@ 2 Dataset Construction
Inspired by Zhuang et al. (2023), we aim to collect data such that the knowledge necessary to answer the questions is absent from LLMs. Therefore, LLMs must consult external resources to answer correctly. Specifically, we mainly collect data from two categories: New world knowledge that is out of the scope of the LLMs’ pre-training corpora. LLMs are static after training and can quickly be outdated due to the ever-changing world (Zhang et al., 2023). To ensure the knowledge is novel to most LLMs, we select 397 QA pairs ranging from 1 October 2023 to 12 January 2024 from RealTimeQA (Kasai et al., 2023). These data comprise weekly quizzes extracted from news websites, encompassing broad topics, including politics, business, and entertainment. In addition, we collect 127 fast-changing questions from FreshQA (Vu et al., 2023), where the answers may change frequently, thereby challenging LLMs’ parametric memorization. Long-tail knowledge that is rarely learned during pre-training. Previous studies (Kandpal et al., 2023) have shown that LLMs struggle to learn less common knowledge and perform poorly without the help of retrieval. Following Asai et al. (2023b), we use the long-tail subset of PopQA (Mallen et al., 2023), which consists of 1,399 rare entity queries with monthly Wikipedia page views below 100, and the test split of unfiltered TriviaQA (Joshi et al., 2017), which has 7,313 factual QA pairs. Lastly, we collect 100 personal agenda questions from ToolQA (Zhuang et al., 2023), which are synthesized with virtual names and events.
@@
As discussed in section˜1, to ensure the questions cannot be answered without external knowledge, we conduct strict filtering. To save manual work, we prompt GPT-4 for answers in a closed-book QA setting without access to external knowledge (see prompt template Fig.LABEL:fig_prompt_no_retrieval in Appendix). Then, we calculate the token-level F1 scores (Rajpurkar et al., 2016) and remove questions that have shared tokens between the prediction and the ground truth, i.e., only keep questions with . Our rationale is that if state-of-the-art GPT-4 cannot answer correctly without retrieval, weaker LLMs are also highly likely to fail. Finally, after filtering, we have obtained 1,271 out of 9,336 questions, covering new world and long-tail knowledge and spanning diverse topics.
@@ 3.1 Standard & Adaptive RAG for QA
| Baselines | Adaptive Retrieval | Always Retrieval | |
@@ 3.2 Experiment Setup
For Self-RAG, we set the retrieval threshold . Lower thresholds encourage more frequent retrieval, while None means the model itself decides when to retrieve by generating the specific [Retrieval] token. Since the quality of the retrieved documents is not the focus of this paper, for long-tail knowledge questions, we use the off-the-shelf Contriever (Izacard et al., 2022) and author-provided top-5 documents extracted from Wikipedia where possible. For questions from ToolQA, we use the author-provided vector database for retrieval of synthesized agendas. Otherwise, we use top-5 documents returned by Google search for new world knowledge questions. We ask LLMs to respond "I don’t know" if they cannot answer the question. More implementation details and prompt templates are in Appendix LABEL:append_implementation_details and LABEL:appendix_prompt.
@@ 4.1 Method
Based on our findings in section˜3.4, we propose Time-Aware Adaptive REtrieval via ICL (TA-ARE), a simple yet effective method to improve ARAG without calibration or additional training. Given that new world knowledge questions often contain time-sensitive information (e.g., "last week", "recent"), we include "Today is current_date()" in the instruction to enhance models’ awareness of time. For long-tail knowledge, we use SimCSE (Gao et al., 2021) to select top-2 semantically closest long-tail questions answered incorrectly from the discarded set in section˜2, denoted as [Yes] demonstrations. For [No] demonstrations, we manually create another two questions, ensuring no extra information is required for most LLMs to answer (e.g., What is the capital of France? ).
@@ 4.3 Ablation Study
The ablation studies in Table 3 validate the effectiveness of TA-ARE: time awareness and relevant in-context demonstrations help LLMs decide the necessity of retrieval for new world and long-tail questions. Table LABEL:tab_fewshot_no_time_results and Table LABEL:tab_no_fewshot_time_results in the Appendix further show the fine-grained performance of each model. We further evaluate the number of in-context demonstrations in Fig.4, showing that 4 demonstrations, comprising 2 [Yes] and 2 [No] examples, have the best performance.
@@ 5 Conclusion
We mainly collect data from existing data sources and use GPT-4 for filtering out answerable questions. While we have done preliminary human checking in Appendix LABEL:append_no_retrieval, it is possible that some questions in the dataset do not require additional information for LLMs to answer. Future work could develop advanced algorithms to do more efficient and rigour filtering.
@@ Appendix A Appendix
A.1 Details of RetrievalQA Constructions
In this work, we collect data that are guaranteed cannot be answered without external information. The dataset statistics are shown in Table 4. The examples of data instances are in Table 1. We collect data from the following sources:
@@ A.1 Details of RetrievalQA Constructions
| Category | Data Source | # Original | # After Filtering | # Avg. Q Tokens | # Avg. Ans Tokens | # Avg. Doc Tokens (Top-5) | ||
| New world knowledge |
| Long-tail knowledge |
@@ A.2 Model Details
We evaluate strong instruction-tuned models with a varying scale of model size: TinyLlama (1.1B) (Zhang et al., 2024), Phi-2222We acknowledge that Phi-2 has not been instruction fine-tuned; however, we find it performs decently well in understanding instructions. (2.7B) (Gunasekar et al., 2023; Li et al., 2023), Llama-2 (7B) (Touvron et al., 2023), GPT-3.5 (OpenAI, 2022), and GPT-4 (OpenAI, 2023). We also use Self-RAG (7B, Asai et al., 2023b), which is fine-tuned based on Llama-2 7B using instruction-following corpora with interleaving passages and reflection tokens. We download the models from HuggingFace333https://huggingface.co/models. The model details, including downloading URLs, model size, and release date, can be found in Table LABEL:tab_model_list.

### v2

@@ Abstract
Adaptive retrieval-augmented generation (ARAG) aims to dynamically determine the necessity of retrieval for queries instead of retrieving indiscriminately to enhance the efficiency and relevance of the sourced information. However, previous works largely overlook the evaluation of ARAG approaches, leading to their effectiveness being understudied. This work presents a benchmark, RetrievalQA, comprising 1,271 short-form questions covering new world and long-tail knowledge. The knowledge necessary to answer the questions is absent from LLMs; therefore, external information must be retrieved to answer correctly. This makes RetrievalQA a suitable testbed to evaluate existing ARAG methods. We observe that calibration-based methods heavily rely on threshold tuning, while vanilla prompting is inadequate for guiding LLMs to make reliable retrieval decisions. Based on our findings, we propose Time-Aware Adaptive REtrieval (TA-ARE), a simple yet effective method that helps LLMs assess the necessity of retrieval without calibration or additional training111The dataset and code are available at https://github.com/hyintell/RetrievalQA.
@@ 1 Introduction
Retrieval-augmented generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023) that augments large language models (LLMs) with retrieval of relevant information has become increasingly popular in knowledge-intensive tasks, including open-domain question-answering (QA) (Zhang et al., 2023b; Kasai et al., 2023; Cui et al., 2023; Zhang et al., 2023a). However, standard RAG methods conduct retrieval indiscriminately, irrespective of the input query, which may result in suboptimal task performance and increased inference costs (Gao et al., 2024). On one hand, LLMs encode vast knowledge in parameters through large-scale pre-training, enabling them to effortlessly handle straightforward queries without retrieval (Mallen et al., 2023). On the other hand, the retrieved context may contain noise and irrelevant information, and augmenting noisy context can potentially distract LLMs, thereby impeding task performance (Shi et al., 2023).
[…]
@@
To fill this gap, we create RetrievalQA, a short-form QA dataset, covering new world and long-tail knowledge and spanning diverse topics. We ensure the knowledge necessary to answer the questions is absent from LLMs. Therefore, LLMs must truthfully decide whether to retrieve to be able to answer the questions correctly. RetrievalQA enables us to evaluate the effectiveness of ARAG approaches, an aspect predominantly overlooked in prior studies and recent RAG evaluation systems (Chen et al., 2023; Saad-Falcon et al., 2023; Es et al., 2023), which focus only on task performance, the relevance of retrieval context or the faithfulness of answers.
@@ 2 Dataset Construction
Inspired by Zhuang et al. (2023), we aim to collect data such that the knowledge necessary to answer the questions is absent from LLMs. Therefore, LLMs must consult external resources to answer correctly. Specifically, we mainly collect data from two categories: New world knowledge that is out of the scope of the LLMs’ pre-training corpora. LLMs are static after training and can quickly be outdated due to the ever-changing world (Zhang et al., 2023b). To ensure the knowledge is novel to most LLMs, we select 397 QA pairs ranging from 1 October 2023 to 12 January 2024 from RealTimeQA (Kasai et al., 2023). These data comprise weekly quizzes extracted from news websites, encompassing broad topics, including politics, business, and entertainment. In addition, we collect 127 fast-changing questions from FreshQA (Vu et al., 2023), where the answers may change frequently, thereby challenging LLMs’ parametric memorization. Long-tail knowledge that is rarely learned during pre-training. Previous studies (Kandpal et al., 2023) have shown that LLMs struggle to learn less common knowledge and perform poorly without the help of retrieval. Following Asai et al. (2023b), we use the long-tail subset of PopQA (Mallen et al., 2023), which consists of 1,399 rare entity queries with monthly Wikipedia page views below 100, and the test split of unfiltered TriviaQA (Joshi et al., 2017), which has 7,313 factual QA pairs. Lastly, we collect 100 personal agenda questions from ToolQA (Zhuang et al., 2023), which are synthesized with virtual names and events.
@@
As discussed in section 1, to ensure the questions cannot be answered without external knowledge, we conduct strict filtering. To save manual work, we prompt GPT-4 for answers in a closed-book QA setting without access to external knowledge (see prompt template Fig.7 in Appendix). Then, we calculate the token-level F1 scores (Rajpurkar et al., 2016) and remove questions that have shared tokens between the prediction and the ground truth, i.e., only keep questions with . Our rationale is that if state-of-the-art GPT-4 cannot answer correctly without retrieval, weaker LLMs are also highly likely to fail. Finally, after filtering, we have obtained 1,271 out of 9,336 questions, covering new world and long-tail knowledge and spanning diverse topics. To avoid potential bias in the evaluation towards methods that retrieve more often, we additionally collect 1,514 questions that can be answered using GPT-2’s parametric knowledge from the discard set. More details are in the Appendix A.2.
@@ 3.1 Standard & Adaptive RAG for QA
| Baselines (1,271) | Adaptive Retrieval | Always Retrieval | |
@@ 3.2 Experiment Setup
For Self-RAG, we set the retrieval threshold . Lower thresholds encourage more frequent retrieval, while None means the model itself decides when to retrieve by generating the specific [Retrieval] token. Since the quality of the retrieved documents is not the focus of this paper, for long-tail knowledge questions, we use the off-the-shelf Contriever (Izacard et al., 2022) and author-provided top-5 documents extracted from Wikipedia where possible. For questions from ToolQA, we use the author-provided vector database for retrieval of synthesized agendas. Otherwise, we use top-5 documents returned by Google search for new world knowledge questions. To reduce API costs, for GPT-4, we randomly select 50 data instances from each source for evaluation, resulting in 250 questions. We ask LLMs to respond "I don’t know" if they cannot answer the question. Due to page limitation, we primarily evaluate the 1,271 questions that need retrieval in the main body and provide the overall results in the Appendix A.6. More implementation details and prompt templates are in Appendix A.4 and A.7.
@@ 4.1 Method
Based on our findings in section 3.4, we propose Time-Aware Adaptive REtrieval via ICL (TA-ARE), a simple yet effective method to improve ARAG without calibration or additional training. Given that new world knowledge questions often contain time-sensitive information (e.g., "last week", "recent"), we include "Today is current_date()" in the instruction to enhance models’ awareness of time. For long-tail knowledge, we use SimCSE (Gao et al., 2021) to select top-2 semantically closest long-tail questions answered incorrectly from the discarded set in section 2, denoted as [Yes] demonstrations. For [No] demonstrations, we manually create another two questions, ensuring no extra information is required for most LLMs to answer (e.g., What is the capital of France? ).
@@ 4.3 Ablation Study
The ablation studies in Table 2 validate the effectiveness of TA-ARE: time awareness and relevant in-context demonstrations help LLMs decide the necessity of retrieval for new world and long-tail questions. Table 8 and Table 9 in the Appendix further show the fine-grained performance of each model. We further evaluate the number of in-context demonstrations in Fig.4, showing that 4 demonstrations, comprising 2 [Yes] and 2 [No] examples, have the best performance.
@@ 5 Conclusion
We mainly collect data from existing data sources and use GPT-4 for filtering out answerable questions. While we have done preliminary human checking in Appendix A.5, it is possible that some questions in the dataset do not require additional information for LLMs to answer. Future work could develop advanced algorithms to do more efficient and rigour filtering.
@@ Appendix A Appendix
A.1 Data Sources
In this work, we collect data from the following sources:
@@ A.1 Data Sources
A.2 Details of RetrievalQA Construction
In this section, we discuss the details of our dataset construction. Our goal is to evaluate adaptive RAG (ARAG) methods and see how good they are at deciding when to retrieve. Therefore, we need the ground truth labels for each question’s retrieval necessity. Ultimately, there are three kinds of questions here:
Case 1: for all LLMs, questions that can be answered using only their parametric knowledge
Case 2: for all LLMs, questions that can not be answered using only the parametric knowledge, therefore requiring external retrieval
Case 3: questions that can be answered with their parametric knowledge for some models but can not be answered for some other models
We do not consider Case 3 because those questions cannot fairly measure whether retrieval is required for different LLMs. For Case 1, it is not trivia to collect questions that can be answered only using the parametric knowledge of LLMs. This is because different LLMs have different levels of pre-trained knowledge, and it is hard to measure (Petroni et al., 2019; Kadavath et al., 2022). For example, given a question, GPT-3.5 may fail to answer and need the help of external knowledge, while Llama-2 may answer correctly using its own knowledge because it has seen the question in the training data. The pre-training corpora are sometimes unavailable, especially for proprietary models, and we can not guarantee that the collected questions can be 100% answered with their own knowledge, as shown in the table below.
However, different from Cases 1 and 3, for Case 2, theoretically, it is possible to collect data that guarantees the knowledge to answer the questions is not present in the models. For instance, new world knowledge occurred after model training and long-tail knowledge that did not (or rarely) appear in the training corpora. Therefore, we collect 1,271 questions (Case 2) that are guaranteed cannot be answered without external information. The data collection process is detailed in section 2. The dataset statistics are shown in Table 3. The examples of data instances are in Table 11.
To avoid potential bias in the evaluation towards methods that retrieve more often, we additionally collect 1,514 questions (Case 1) that can be answered using GPT-2’s parametric knowledge from the discard set. Specifically, we use GPT-2 (small, 124M) in the zero-shot closed-book QA setting to evaluate the discard set. We only keep questions that can be answered using GPT-2’s parametric knowledge (when the loose match score = 1), assuming that larger and stronger LLMs are also highly likely to succeed if small and weak GPT-2 can answer correctly without retrieval. We also use the entire PopQA dataset (the rest of the long-tail split), which has 12,883 data instances that are more common on the web. We found that GPT-2 cannot answer any new-world questions from the discard set, which is reasonable.
| Category | Data Source | # Original | # After Filtering | # Avg. Q Tokens | # Avg. Ans Tokens | # Avg. Doc Tokens (Top-5) |
| New world knowledge | RealTimeQA (Kasai et al., 2023) | |||||
| FreshQA (Vu et al., 2023) | ||||||
| Long-tail knowledge | ToolQA (Zhuang et al., 2023) | |||||
| PopQA (Mallen et al., 2023) | ||||||
| TriviaQA (Joshi et al., 2017) | ||||||
@@ A.3 Model Details
We evaluate strong instruction-tuned models with a varying scale of model size: TinyLlama (1.1B) (Zhang et al., 2024), Phi-2222We acknowledge that Phi-2 has not been instruction fine-tuned; however, we find it performs decently well in understanding instructions. (2.7B) (Gunasekar et al., 2023; Li et al., 2023), Llama-2 (7B) (Touvron et al., 2023), GPT-3.5 (OpenAI, 2022), and GPT-4 (OpenAI, 2023). We also use Self-RAG (7B, Asai et al., 2023b), which is fine-tuned based on Llama-2 7B using instruction-following corpora with interleaving passages and reflection tokens. We download the models from HuggingFace333https://huggingface.co/models. The model details, including downloading URLs, model size, and release date, can be found in Table A.3.
lCc[]
Model Name Model Size Release
TinyLlama/TinyLlama-1.1B-Chat-v1.0 1.1B Dec 2023
microsoft/phi-2 2.7B Dec 2023
meta-llama/Llama-2-7b-chat-hf 7B Jul 2023
selfrag/selfrag_llama2_7b 7B Oct 2023
gpt-3.5-turbo – Nov 2022
gpt-4-turbo-preview – Mar 2023
| Parameters | Values |
| temperature | |
| top_p | |
| max_tokens | |
| Retrieved docs | |
| Threshold (Self-RAG) | |
| # demonstrations (TA-ARE) | |
| Eval metric | match/retrieval accuracy |
A.4 Implementation Details
For fair comparisons, we use the same setting following Self-RAG for all experiments. The detailed hyperparameters are summarized in Table 5.
For Self-RAG, we set the retrieval threshold . Lower thresholds encourage more frequent retrieval, while None means the model itself decides when to retrieve by generating the specific [Retrieval] token. Since the quality of the retrieved documents is not the focus of this paper, we use the off-the-shelf Contriever (Izacard et al., 2022) and author-provided top-5 documents extracted from Wikipedia where possible for long-tail knowledge questions. For questions from ToolQA, we use the author-provided vector database for retrieval of synthesized agendas. Otherwise, we use top-5 documents returned by Google search444We use SerpApi for Google search. for new world knowledge questions. To reduce API costs, for GPT-4, we randomly select 50 data instances from each source for evaluation, resulting in 250 questions. We ask LLMs to respond "I don’t know" if they cannot answer the question.
For instruction-tuned LMs, we use the official system prompt or instruction format used during training if publicly available. We use vLLM (Kwon et al., 2023) for accelerated inference.
A.5 Sanity Check: Baselines Without Retrieval
We perform a sanity check on our RetrievalQA (questions need retrieval) using a simple QA template (Fig.7) without retrieval. As shown in Table A.5 and Fig.1, all models achieve very poor match and F1 scores on RetrievalQA, indicating that it is extremely hard for models to answer the questions without consulting external resources.
We notice that TinyLlama, Phi-2, and Self-RAG have slightly better performance than larger models. Considering that these models were trained recently (as shown in Table A.3), they might have learned some new knowledge and can answer some questions correctly. Additionally, we conducted human checking on the questions answered correctly and found that some questions were mismarked due to multiple possible ground truths. For example, for the question: "Where will NeurIPS be located this year (2024)?", the model answers: "NeurIPS will be held in Montreal, Canada.", and the ground truth is an array of ["Vancouver, Canada", "Vancouver", "Canada"]. Since the model prediction contains Canada, this answer was marked correct. However, LLMs themselves still do not truly know the answer. The outdated knowledge stored in their parameters makes them hallucinate. Since these questions only take a tiny portion of the entire dataset (as an example shown in Fig.1, the tiny red line from Retrieval=No to Prediction=Correct), and early-trained models such as Llama-2 and GPT-3/4 perform worse, we still keep them in our dataset.
lCC
Model Match F1
TinyLlama (1.1B) 4.2 1.3
Phi-2 (2.7B) 7.2 3.9
Llama-2 (7B) 2.0 0.7
Self-RAG (7B) 6.0 1.5
GPT-3.5 1.2 1.0
GPT-4* 2.4 2.3
lC
Model Match
TinyLlama (1.1B) 88.1
Phi-2 (2.7B) 87.7
Llama-2 (7B) 89.8
Self-RAG (7B) 88.2
GPT-3.5 91.1
GPT-4* 88.4
| Baselines | Adaptive Retrieval | |
| Retrieval | Match | |
| TinyLlama (1.1B) | ||
| Phi-2 (2.7B) | ||
| Llama-2 (7B) | ||
| GPT-3.5 | ||
| GPT-4 | ||
| Average gain | ||
| Baselines | Adaptive Retrieval | |
| Retrieval | Match | |
| TinyLlama (1.1B) | ||
| Phi-2 (2.7B) | ||
| Llama-2 (7B) | ||
| GPT-3.5 | ||
| GPT-4 | ||
| Average gain | ||
We also run the sanity check on the 1,514 questions that do not need retrieval. As shown in Table A.5, even strong models like GPT-3.5 and GPT-4 can not reach 100% match accuracy using their parametric knowledge.
| Baselines (2,785) | No Retrieval | Adaptive Retrieval | Always Retrieval | ||||
| Match | Match | Retrieval Acc | Precision | Recall | F1 | Match | |
| Calibration-based | |||||||
| Self-RAG (7B) | |||||||
| Model-based | |||||||
| Vanilla section 3 | |||||||
| TinyLlama (1.1B) | |||||||
| Phi-2 (2.7B) | |||||||
| Llama-2 (7B) | |||||||
| GPT-3.5 | |||||||
| GPT-4* | |||||||
[…]
| TinyLlama (1.1B) | |||||||
| Phi-2 (2.7B) | |||||||
| Llama-2 (7B) | |||||||
| GPT-3.5 | |||||||
| GPT-4* | |||||||
| Average gain | – | – | |||||
A.6 Overall Results
In Table 1, we only evaluate 1,271 questions that need retrieval. In this section, we evaluate the entire 2,785 data, with 1,271 labelled as required retrieval and 1,514 labelled as do not require retrieval. Besides retrieval accuracy, we also report retrieval macro precision, recall, and F1. Table 10 shows the overall results. Using questions that do not need retrieval and questions that need retrieval, we have a comprehensive evaluation of ARAG methods.
| Category | Data Source | Question | Answer |
| New world knowledge | RealTimeQA (Kasai et al., 2023) | Which 2024 Republican presidential contender announced that he is ending his campaign? | Former Texas Rep. Will Hurd |
| FreshQA (Vu et al., 2023) | What is the latest highest-grossing movie of the week at the Box office? | Mean Girls | |
| Long-tail knowledge | ToolQA (Zhuang et al., 2023) | What time did Grace attend Broadway Show on 2022/02/17? | 8:00 PM |
| PopQA (Mallen et al., 2023) | What is Henry Feilden’s occupation? | politician | |
| TriviaQA (Joshi et al., 2017) | Which bird, that breeds in northern Europe in pine and beech forests, has a chestnut brown back, grey head, dark tail, buff breast and a striped black throat? | fieldfare |